OpenStudy (anonymous):

Could someone help me learn how to do this?? I will give a medal and fan to whoever help me learn the best! Find the exact value of the expression 4 sqrt 13 + sqrt 6 - 7 sqrt 13

OpenStudy (reemii):

\(\sqrt{2}+\sqrt2 = 2\sqrt{2}\) right? similarly, \(4\sqrt{13} - 7\sqrt{13} = -3\sqrt{13}\). Can you "put together" \(\sqrt{13}\) and \(\sqrt{6}\) ?

OpenStudy (anonymous):

@reemii i'm not sure

OpenStudy (reemii):

you can't. the numbers under the root must be equal. the answer is a sum of two terms (you put the \(\sqrt{13}\) together as i wrote above alr.

OpenStudy (anonymous):

So \[-3\sqrt{13} + \sqrt{16}\] ???

OpenStudy (anonymous):

@reemii

OpenStudy (reemii):

yes (but it's 6, not 16)

OpenStudy (anonymous):

Oops, my bad, thank you!